The 9-month Libor rate is 4%. The 12-month Libor rate is 5%. An investor creating a homemade forward forward could earn what annualized interest rate on a 3-month $1,000,000 deposit made in 9 months?

I'm pretty confident it is:

1,000,000 / 1 + .04 (3/4) = $970,874

$970,874 x (1 + .05 (4/4)) = $1,019,418

then this is where I'm unsure... is it:

$1,019,418 - $970,874 = $48,544

$48,544 / $1,000,000 = .0485 = 4.85%

Please show work if this isn't right
Answer #1

Hi,

You are correct for 1st step.

1,000,000 / 1 + .04 (3/4) = $970,874

$970,874 x (1 + .05 (4/4)) = $1,019,418

Now read the question once again and check what you really want as a solution.

You need interest for 3 months 9 months from now.

Investment amount at the end of 9 months = $ 1,000,000

Value of Investment at the end of 12 months = $ 1,019,418

Therefore interest for 3 months = 19,418

Therefore interest rate for 3 months = 19,418 / 1,000,000 x 12/3 = 7.767% p.a.